To avoid detection at customs, a traveler places 6 narcotic tablets in a bottle containing 9 vitamin tablets that are similar in appearance. If the customs official selects 3 of the tablets at random for analysis, what is the probability that the traveler will be arrested for illegal possession of narcotics?


1
Expert's answer
2021-07-14T12:57:40-0400

"N="Total number of tablets placed in the bottle"=9+6=15"

"m=" Total number of narcotic drugs in the bottle"=6"

"n=" Total number of tablets selected randomly by the customs official"=3"


Let "x="The number of narcotic tablet(s) selected in the sample.

Since the selections are made without replacement hence, the distribution of "x" would be:

"x" "Hypergeometric (N=15, m=6, n=3)"

The proof of "x" becomes:

"P(x=x)=\frac{\begin{pmatrix}m\\ x\end{pmatrix}\begin{pmatrix}N-m\\ \:n-x\end{pmatrix}}{\begin{pmatrix}N\\ \:n\end{pmatrix}}\:\:x=max\left(0,\:n+m-N\right),\:\:min\left(n,\:m\right)"

"Hence,\:max\left(0,\:n+m-N\right)\:and\:min\left(n,m\right)"

"=max\left(0,\:3+6-15\right)\:and\:min\left(3,\:6\right)"

"=max\left(0,\:-6\right)=0\:and\:min=3"


The proof of "x" becomes:

"\:P\left(x=x\right)=\frac{\begin{pmatrix}6\\ x\end{pmatrix}\begin{pmatrix}9\\ 3-x\end{pmatrix}}{\begin{pmatrix}15\\ 3\end{pmatrix}};\:x=0,\:1,\:2,\:3"

Therefore we need,

"P\left(The\:travellers\:will\:be\:arrested\:for\:illegal\:possesion\:of\:narcortics\right)="

"P\left(There\:will\:be\:at\,least\:one\:narcotic\:tablet\:in\:the\:sample\right)="

"P\left(x\ge 1\right)="

"1-P\left(x<1\right)" "\left(Law\:of\:complement\:events\right)="

"1-P\left(x=0\right)"

"=1-\frac{\begin{pmatrix}6\\ 0\end{pmatrix}\begin{pmatrix}9\\ 3\end{pmatrix}}{\begin{pmatrix}15\\ 3\end{pmatrix}}\:\:\left(By\:the\:proof\:of\:x\right)" "\approx" "0.8154"